OpenStudy (anonymous):

PLEASE HELP! WILL FAN AND MEDAL! Which options could be the next step in simplifying this expression? 27 + (3 + 41) Choose exactly two answers that are correct. A. 27 + 44 (use order of operations) B. 30 + 68 (use the distributive property) C. 30 + 41 (use the associative property) D. 68 + 30 (use the commutative property)

OpenStudy (danjs):

the parenthesis show that you should do that calculation first terms added together can be done in any order, so you can group any terms you want together and not change the result

OpenStudy (danjs):

27 + ( 3 + 41) = (27 + 3) + 41 = (27 + 41 ) + 3

OpenStudy (anonymous):

So A and D?

OpenStudy (danjs):

Not D, that sums up different than the problem

OpenStudy (anonymous):

So B?

OpenStudy (danjs):

do you see B) 30 + 68 = 98 the problem, 27 + 3 + 41 = 71 not the same

OpenStudy (anonymous):

it would be A.

OpenStudy (danjs):

A applies the order of operations by adding the parenthesis first C uses the associative property by changing where the parenthesis are (27+3) + 41

OpenStudy (ahvei):

A and C are the answers

OpenStudy (ahvei):

hope this helps